# Effective resource allocation to combat invasions of the spotted lanternfly (Lycorma delicatula) and similar pests

### v1

@@ Discussion
$\frac{\partial\lambda^{x}}{\partial p_{i}}<\frac{\partial\lambda^{x}}{\partial p_{j}},$
where the switch in inequality direction results from the fact that the partial derivatives are strictly negative. We also note that, when formulated in more general terms, this is a well-known result typically established via linear programming [39].
@@
$\frac{dp^{d}}{d\epsilon}\bigg{|}_{\epsilon=0}=r_{i}e^{-r_{i}\cdot 0}=r_{i}$
@@
$\frac{dp^{p}}{d\epsilon}\bigg{|}_{\epsilon=0}=1-k_{i},$
@@
$\lambda^{x}(\hat{\epsilon})=\lambda_{0}\prod_{i=1}^{4}\sigma_{i}(\epsilon_{i})$
and note that the derivative of $\lambda^{x}(\hat{\epsilon})$ with respect to $\epsilon_{i}$ is given by
$\frac{\partial\lambda^{x}}{\partial\epsilon_{i}}=\lambda_{0}r_{i}^{2}e^{-r_{i}\epsilon_{i}}\prod_{j\neq i}\sigma_{j}(\epsilon_{j}).$
@@
$\frac{\partial\lambda^{x}}{\partial\epsilon_{i}}=\frac{\partial\lambda^{x}}{\partial\epsilon_{i+1}}\bigg{|}_{\epsilon_{i+1}=0}$
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\prod_{j\neq i}\sigma_{j}(\epsilon_{j})=r_{i+1}^{2}\prod_{j\neq i+1}\sigma_{j}(\epsilon_{j})$
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\sigma_{i+1}(0)=r_{i+1}^{2}\sigma_{i}$
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\left(1-r_{i+1}(1-e^{-r_{i+1}0})\right)=r_{i+1}^{2}\left(1-r_{i}(1-e^{-r_{i}\epsilon_{i}})\right)$
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}=r_{i+1}^{2}\left(1-r_{i}(1-e^{-r_{i}\epsilon_{i}})\right)$
$\frac{r_{i}^{2}}{r_{i+1}^{2}}=e^{r_{i}\epsilon_{i}}\left(1-r_{i}\right)+r_{i}$
$e^{r_{i}\epsilon_{i}}\left(1-r_{i}\right)=\frac{r_{i}^{2}}{r_{i+1}^{2}}-r_{i}$
$e^{r_{i}\epsilon_{i}}=\frac{r_{i}^{2}-r_{i}r_{i+1}^{2}}{r_{i+1}^{2}(1-r_{i})}=\frac{r_{i}(r_{i}-r_{i+1}^{2})}{r_{i+1}^{2}(1-r_{i})}$
$\epsilon_{i}=\frac{1}{r_{i}}\ln\frac{r_{i}(r_{i}-r_{i+1}^{2})}{r_{i+1}^{2}(1-r_{i})}.$
@@
To perform the constrained global numerical optimization we used the Mathematica function Minimize [38]. For the example in Figure 1 the following command was used Minimize[{5.47*(1 - 0.9*(1 - Exp[-0.9 x]))*(1 - 0.85 * (1 - Exp[-0.85 y]))*(1 - 0.65*(1 - Exp[-0.65 z]))*(1 - 0.5*(1 - Exp[-0.5 w])), x + y + z + w == 10 && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}] where $x$, $y$, $z$ and $w$ are placeholders for $\epsilon_{1}$, $\epsilon_{2}$, $\epsilon_{3}$ and $\epsilon_{4}$, respectively. The exact values obtained were: final minimum growth rate 0.061416, $\epsilon_{1}=3.78794$, $\epsilon_{2}=3.37848$, $\epsilon_{3}=2.12919$, $\epsilon_{4}=0.704396$. This command can be modified to analyze situations with different budgets $B$ and stage-specific growth rates $r_{i}$ (or $k_{i}$) by substituting these into this command Minimize[{5.47*(1 - r1 * (1 - Exp[-r1 x]))*(1 - r2 * (1 - Exp[-r2 y]))*(1 - r3*(1 - Exp[-r3 z]))*(1 - r4*(1 - Exp[-r4 w])), x + y + z + w == B && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}]. It can also be modified to work with entirely different $p_{i}=p_{i}(\epsilon_{i})$ functions by replacing the current ones, i.e. Minimize[{5.47*(1 - p1(x))*(1 - p2(y))*(1 - p3(z))*(1 - p4(w)), x + y + z + w == B && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}]. More stages can also be included by adding variables and inserting additional factors into the product that correspond to those stages.

### v2

@@ Discussion
$\frac{\partial\lambda^{x}}{\partial p_{i}}<\frac{\partial\lambda^{x}}{\partial p_{j}},$ (19)
where the switch in inequality direction results from the fact that the partial derivatives are strictly negative. We also note that, when formulated in more general terms, this is a well-known result typically established via linear programming (Dantzig1963, ).
Example numerical illustration of Result 2
With $\lambda^{x}(\bar{p})=\lambda_{0}\prod_{i=1}^{4}(1-p_{i}(1-k_{i}))$, the partial derivatives are $\frac{\partial\lambda^{x}}{\partial p_{i}}=-\lambda_{0}\\ (1-k_{i})\prod_{j\neq i}(1-p_{j}(1-k_{j})).$ At $\bar{p}=\bar{0}$ this reduces to $\frac{\partial\lambda^{x}}{\partial p_{i}}\big{|}_{\bar{p}=\mathbf{0}}=-\lambda_{0}(1-k_{i})$, so the stage with the smallest $k_{i}$ yields the largest marginal reduction in $\lambda^{x}$. Using the values corresponding to the example in Figure 1, i.e. $k_{1}=0.10$, $k_{2}=0.15$, $k_{3}=0.35$, $k_{4}=0.50$ gives
$\frac{\partial\lambda^{x}}{\partial p_{1}}=-0.90\lambda_{0},\quad\frac{\partial\lambda^{x}}{\partial p_{2}}=-0.85\lambda_{0},\quad\frac{\partial\lambda^{x}}{\partial p_{3}}=-0.65\lambda_{0},\quad\frac{\partial\lambda^{x}}{\partial p_{4}}=-0.50\lambda_{0},$
hence $|\partial\lambda^{x}/\partial p_{1}|>|\partial\lambda^{x}/\partial p_{2}|>|\partial\lambda^{x}/\partial p_{3}|>|\partial\lambda^{x}/\partial p_{4}|$, which matches Result 2: prioritize the smaller $k_{i}$ first. The same ordering holds for nonzero $\bar{p}$ whenever $p_{i}>p_{j}\frac{1-k_{j}}{1-k_{i}}$, as stated in the result.
@@
$\frac{dp^{d}}{d\epsilon}\bigg{|}_{\epsilon=0}=r_{i}e^{-r_{i}\cdot 0}=r_{i}$ (20)
@@
$\frac{dp^{p}}{d\epsilon}\bigg{|}_{\epsilon=0}=1-k_{i},$ (21)
@@
In empirical applications, however, $r_{i}$ need not be set equal to $1-k_{i}$. Instead, $r_{i}$ can be estimated directly from effort–coverage data by fitting the function $p_{i}(\epsilon_{i})=1-e^{-r_{i}\epsilon_{i}}$. For a single observation $(\epsilon_{i},p_{i})$ we can find $r_{i}$ through $r_{i}=-\ln(1-p_{i})/\epsilon_{i}$, and for multiple observations $r_{i}$ can be obtained by least-squares regression. This provides a direct link between field data (e.g. trap density, spray volume, or labor effort ($\epsilon_{i}$) and proportion affected ($p_{i}$)) and the parameter $r_{i}$.
@@
$\lambda^{x}(\hat{\epsilon})=\lambda_{0}\prod_{i=1}^{4}\sigma_{i}(\epsilon_{i})$ (22)
[…]
$\frac{\partial\lambda^{x}}{\partial\epsilon_{i}}=\lambda_{0}r_{i}^{2}e^{-r_{i}\epsilon_{i}}\prod_{j\neq i}\sigma_{j}(\epsilon_{j}).$ (23)
@@
$\frac{\partial\lambda^{x}}{\partial\epsilon_{i}}=\frac{\partial\lambda^{x}}{\partial\epsilon_{i+1}}\bigg{|}_{\epsilon_{i+1}=0},$ (24)
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\prod_{j\neq i}\sigma_{j}(\epsilon_{j})=r_{i+1}^{2}\prod_{j\neq i+1}\sigma_{j}(\epsilon_{j}),$ (25)
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\sigma_{i+1}(0)=r_{i+1}^{2}\sigma_{i},$ (26)
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}\left(1-r_{i+1}(1-e^{-r_{i+1}0})\right)=r_{i+1}^{2}\left(1-r_{i}(1-e^{-r_{i}\epsilon_{i}})\right),$ (27)
$r_{i}^{2}e^{-r_{i}\epsilon_{i}}=r_{i+1}^{2}\left(1-r_{i}(1-e^{-r_{i}\epsilon_{i}})\right),$ (28)
$\frac{r_{i}^{2}}{r_{i+1}^{2}}=e^{r_{i}\epsilon_{i}}\left(1-r_{i}\right)+r_{i},$ (29)
$e^{r_{i}\epsilon_{i}}\left(1-r_{i}\right)=\frac{r_{i}^{2}}{r_{i+1}^{2}}-r_{i},$ (30)
$e^{r_{i}\epsilon_{i}}=\frac{r_{i}^{2}-r_{i}r_{i+1}^{2}}{r_{i+1}^{2}(1-r_{i})}=\frac{r_{i}(r_{i}-r_{i+1}^{2})}{r_{i+1}^{2}(1-r_{i})},$ (31)
@@
To perform the constrained global numerical optimization we used the Mathematica function Minimize (Mathematica, ). For the example in Figure 1 the following command was used Minimize[{5.47*(1 - 0.9*(1 - Exp[-0.9 x]))*(1 - 0.85 * (1 - Exp[-0.85 y]))*(1 - 0.65*(1 - Exp[-0.65 z]))*(1 - 0.5*(1 - Exp[-0.5 w])), x + y + z + w == 10 && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}] where $x$, $y$, $z$ and $w$ are placeholders for $\epsilon_{1}$, $\epsilon_{2}$, $\epsilon_{3}$ and $\epsilon_{4}$, respectively. The exact values obtained were: final minimum growth rate 0.061416, $\epsilon_{1}=3.78794$, $\epsilon_{2}=3.37848$, $\epsilon_{3}=2.12919$, $\epsilon_{4}=0.704396$. This command can be modified to analyze situations with different budgets $B$ and stage-specific growth rates $r_{i}$ (or $k_{i}$) by substituting these into this command Minimize[{5.47*(1 - r1 * (1 - Exp[-r1 x]))*(1 - r2 * (1 - Exp[-r2 y]))*(1 - r3*(1 - Exp[-r3 z]))*(1 - r4*(1 - Exp[-r4 w])), x + y + z + w == B && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}]. It can also be modified to work with entirely different $p_{i}=p_{i}(\epsilon_{i})$ functions by replacing the current ones, i.e. Minimize[{5.47*(1 - p1(x))*(1 - p2(y))*(1 - p3(z))*(1 - p4(w)), x + y + z + w == B && x >= 0 && y >= 0 && z >= 0 && w >= 0}, {x, y, z, w}]. More stages can also be included by adding variables and inserting additional factors into the product that correspond to those stages.
Code Availability Statement
All MATLAB code required to replicate the numerical results in this paper can be accessed here https://github.com/danielstrombom/SLFOC.